The temperature rose 15 degrees to 7 degrees. What was the original tempeture?
Answer by funmath(2933)

The temperature rose 15 degrees to 7 degrees. What was the original tempeture?
Let the original temperature be:x
Then
x+15=7
x+15-15=7-15
x=-8 degrees
:
Check:
-8+15=7
7=7 We're right!!!
Happy Caclculating!!!